# KernelDNA: Dynamic Kernel Sharing via Decoupled Naive Adapters

### v1

@@ Abstract
Dynamic convolution enhances model capacity by adaptively combining multiple kernels, yet faces critical trade-offs: prior works either (1) incur significant parameter overhead by scaling kernel numbers linearly, (2) compromise inference speed through complex kernel interactions, or (3) struggle to jointly optimize dynamic attention and static kernels. We also observe that pre-trained Convolutional Neural Networks (CNNs) exhibit inter-layer redundancy akin to that in Large Language Models (LLMs). Specifically, dense convolutional layers can be efficiently replaced by derived “child” layers generated from a shared “parent” convolutional kernel through an adapter.
To address these limitations and implement the weight-sharing mechanism, we propose a lightweight convolution kernel plug-in, named KernelDNA. It decouples kernel adaptation into input-dependent dynamic routing and pre-trained static modulation, ensuring both parameter efficiency and hardware-friendly inference. Unlike existing dynamic convolutions that expand parameters via multi-kernel ensembles, our method leverages cross-layer weight sharing and adapter-based modulation, enabling dynamic kernel specialization without altering the standard convolution structure. This design preserves the native computational efficiency of standard convolutions while enhancing representation power through input-adaptive kernel adjustments. Experiments on image classification and dense prediction tasks demonstrate that KernelDNA achieves state-of-the-art accuracy-efficiency balance among dynamic convolution variants. Our codes are available at https://github.com/haiduo/KernelDNA 111Preprint. Under review..
@@ 1 Introduction
Convolutional Neural Networks [8, 23, 21] have long been built on static convolutional kernels within each layer. While these static kernels enable computationally efficient inference, their fixed, input-agnostic nature inherently limits representation flexibility. Dynamic convolution [28, 4] attempts to address this by adaptively modulating kernels based on input content, but it inevitably scales parameters linearly with the number of kernels, introducing prohibitive memory costs that hinder real-world deployment. Existing model compression techniques, such as pruning [16], quantization [5], and knowledge distillation [6], aim to alleviate parameter growth but require complex multi-stage pipelines: they often depend on fine-tuning pre-trained models with carefully tuned hyperparameters. In contrast, weight-sharing [25, 2] strategies in LLMs demonstrate that parameter redundancy can be exploited for drastic compression without performance loss. Inspired by this, we propose integrating dynamic neural networks’ adaptive capacity with LLM-like weight sharing, unlocking the potential of CNNs through a novel parameter-efficient dynamic modulation framework.
@@
To validate the inherent inter-layer redundancy in conventional CNNs, we visualize the similarity between convolutional kernels across layers using the Linear Centered Kernel Alignment (CKA) index [12]. As shown in Fig. 1, except for a few layers that exhibit lower similarity with others, the majority of layers show high similarity. This observation suggests that compression or transformation can be applied to enable similar layers to share a base layer, thereby reducing the model’s parameters without compromising the architecture. Drawing inspiration from Epigenetics [24]—where phenotypic traits are regulated through non-heritable modifications (e.g., DNA methylation) without altering the underlying genetic code—we design a lightweight adapter module. This module decouples input-dependent dynamic channel attention from kernel-centric spatial modulation, enabling a shared “parent” kernel within a block or stage to generate derived “child” layers via adapter-based transformations. By replacing dense convolutional layers with these adapted child layers, we achieve substantial parameter reduction while preserving performance. Fig. 2 illustrates the distinction between the kernel sharing mechanism in our method and mainstream dynamic convolution approaches.
@@
Previous both paradigms (i.e., Fig. 2 (b) and (c)) share critical limitations: they introduce computationally intensive operations (e.g., dynamic kernel assembly or generation) that degrade inference speed and often cause parameter inflation. In contrast, our kernel-sharing strategy not only reduces parameters but also preserves the efficient inference advantage of traditional convolution by reusing and modulating shared base kernels across layers. Notably, parameter sharing introduces implicit regularization [13], mitigating overfitting risks and even enhancing generalization.
@@
To ensure hardware-friendly efficiency, our adapter integrates three enhanced attention mechanisms: Channel Attention, Filter Attention, and Spatial Attention. These mechanisms allow each convolutional layer to retain distinct characteristics while benefiting from cross-layer (common patterns) captured through shared parameters. Termed KernelDNA, this plug-and-play module functions as an epigenetic “marker”, dynamically tailoring kernel behavior without altering the base parameter structure. It is universally applicable to any standard convolution operator, including regular convolutions, depthwise convolutions, and 11 convolutions.
@@
Extensive experiments on ImageNet-1K and downstream tasks demonstrate that KernelDNA surpasses prior dynamic convolution methods in accuracy while preserving near-original inference speeds (e.g., 7,683 fps vs. ResNet18’s 8,463 fps, 90.7% throughput retention), as shown in Fig. 3. Compared to SOTA variants like ODConv (44.90M params) and CondConv (81.35M params), KernelDNA achieves 1.2–5 parameter reduction (9.24M for ResNet18) with higher accuracy (74.23% vs. 73.97% Top-1). On lightweight models like MobileNetV2, KernelDNA attains 75.51% Top-1 accuracy (vs. KernelWarehouse’s 74.68%) using only 3.14M parameters (vs. KernelWarehouse’s 5.17M), while maintaining 99% of the base model’s throughput (6,260 fps vs. 6,323 fps). These results establish a new Pareto frontier, where KernelDNA uniquely balances parameter efficiency (up to 5 compression), hardware compatibility (near-static latency), and adaptive performance across diverse architectures.
@@ 2 Related Work
Dynamic Convolution: Dynamic convolution allows adaptive feature extraction based on input content to address the limited adaptability of static convolution to input variations. Early methods like CondConv [28] proposes conditionally parameterized convolutions, where multiple static kernels are aggregated via input-dependent routing weights. DY-CNNs [4] replaces simple routing with attention mechanisms. DCD [17] leverages kernel matrix decomposition for dynamic channel fusion, eliminating channel-wise dynamic attention. DD [30] decomposed dynamic filters into spatial and channel-specific components, effectively combining position-wise channel and spatial attentions. Recent advances further expanded dynamic convolution’s scope: ODConv [15] introduces dynamic adjustments to kernel spatial size, input/output channel dimensions, and kernel-wise multiplication operations. While improving performance, its reliance on dynamic kernel computations significantly increases FLOPs, degrading inference speed. SD-Conv [10] reduces parameters and FLOPs by sparsifying dynamic kernels, but at the cost of accuracy degradation and unpredictable latency spikes. The SOTA KernelWarehouse [14] assembles convolutional kernels from a fine-grained, static cell-wise kernel bank using input-dependent attention scores. However, the kernel assembly process incurs substantial latency and throughput overhead, making it impractical for resource-constrained devices. In contrast, our KernelDNA leverages a weight-sharing mechanism to preserve the computational efficiency of standard convolution while reducing parameters.
Weight Sharing: The concept of weight sharing originated in NLP, where ALBERT [13] pioneeres cross-layer weight sharing to prevent parameter growth with network depth. This idea later migrated to vision: MiniViT [29] applies weight multiplexing to compress vision transformers while enhancing prediction performance. Recent advances further refine sharing strategies—MPOBERT [20] decomposes parameter matrices into shared local tensors, Dynamic Layer Tying [7] employes reinforcement learning to dynamically select layer-wise sharing relationships during training, and Basis Sharing [25] achieves parameter reuse across layers via singular value decomposition (SVD) for LLM compression. The SOTA Relaxed Recursive Transformers[1] revisits layer tying integrated with Low-Rank Adaptation (LoRA), pushing the frontier of efficient LLM compression. Despite its success in transformers, weight sharing remains underexplored in convolutional networks. To our knowledge, this work is the first to adapt weight-sharing principles to dynamic CNNs, enabling efficient inference and compression.
Visual Attention Module: Visual attention mechanisms enhance the representational power of models by dynamically adjusting feature weights, becoming a core strategy for optimizing CNNs. SE-Net [11] pioneeres channel-wise attention by using global average pooling (Squeeze) and fully connected layers (Excitation) to learn channel weights, amplifying important channels while suppressing redundant ones. CBAM [27] combines channel and spatial attention, generating weights through global pooling and spatial convolutions to calibrate multi-dimensional features. ECA-Net [26] improves SE-Net by replacing fully connected layers with 1D convolutions, avoiding information loss from dimensionality reduction and adaptively adjusting cross-channel interaction coverage. CGC [19] dynamically adjusts convolutional kernel parameters through a context-gating mechanism, integrating input features and global context to generate gating weights, enhancing the contextual relevance of feature representation. In contrast, our KernelDNA adopts a decoupled dynamic-static design, where dynamic Channel Attention operates in parallel with static Spatial/Filter attention, reducing runtime computational dependencies.
@@ 3 Preliminaries
Standard Convolution. Given an input feature map with channels and spatial dimensions , the standard convolution operation applies a fixed filter to produce an output feature map . At each spatial location , the convolution is computed as:
where denotes the neighborhood around , and is the bias term. The filter is shared across all spatial locations, making the operation computationally efficient but input-agnostic.
Dynamic Filters. Dynamic filters extend standard convolution by generating input-dependent filters for each spatial location . This allows the convolution to adapt to local features:
However, generating requires significant computational overhead, limiting its application to small networks.
Dynamic Convolution. Dynamic convolution improves efficiency by aggregating a set of static kernels using input-dependent attention weights :
where is computed via a lightweight attention mechanism. This approach enhances representational power with minimal computational overhead.
@@ 4.1 Dynamic Kernel Sharing
Prior works attempte to enhance the potential of dynamic convolution by scaling it to finer-grained or larger-scale settings. For instance, the recent SOTA method, KernelWarehouse [14], learns linear combinations of n static fine-grained cell-wise convolutional kernels to explore performance with kernel counts exceeding 100. While it decomposes original kernels into finer-grained components and constructs a shared kernel warehouse, the method dynamically assembles full kernels by weighting cell-level kernels via a contrast-driven attention function based on input features. The reconstructed kernels are then convolved with the input.
Although KernelWarehouse claims to enable cross-layer kernel sharing, it fundamentally follows the traditional kernel bank paradigm—aggregating candidate static kernels via attention-based weighting, albeit at a finer granularity. A detailed implementation is shown in Fig. 4 (a). Despite its improved performance over dense convolutions, KernelWarehouse suffers from a critical flaw: computing attention scores for fine-grained kernels and integrating them into full kernels incur substantial computational overhead. These operations are hardware-unfriendly due to intensive memory access, leading to significantly slower inference speeds compared to standard dense convolutions, thus limiting practical deployment in latency-sensitive scenarios.
In contrast to KernelWarehouse’s fine-grained decomposition and reassembly, our KernelDNA preserves the structure of standard convolutions while introducing layer-wise parameter sharing. Instead of complex computations, KernelDNA employs lightweight adapters to generate layer-specific kernels from shared base parameters, retaining the inference efficiency of standard convolutions. This design enables seamless replacement of traditional convolutions, as depicted in Fig. 4 (b). Where the solid boxes represent static convolutional kernels, which serve as the parent layers in KernelDNA. The dashed boxes denote child layers derived from the parent layers through adapter transformations. These child layers do not introduce additional parameters in storage but are dynamically generated during inference to perform convolution operations.
Furthermore, our adapter integrates both static attention (determined during training) and input-dependent dynamic attention (computed at inference), decoupling purely dynamic attention mechanisms of previous works. This allows flexible adaptation to deployment requirements:
High throughput mode: Disable dynamic attention for minimal latency at a slight accuracy cost.
High accuracy mode: Enable dynamic attention for improved performance with marginal speed trade-offs.
KernelDNA can serve as a plug-and-play replacement for conventional convolution layers, balancing efficiency and adaptability without compromising hardware compatibility.
@@ 4.2 Light-weight Modulation Adapter
To efficiently and effectively transform base (parent) kernels into shared (child) kernels, we design a lightweight adapter that incorporates three attention mechanisms to modulate kernel properties across different dimensions: input channel, output channel, and spatial interactions. These mechanisms are termed Channel Attention, Filter Attention, and Spatial Attention, respectively. As illustrated in Fig. 5, Channel Attention is dynamically computed based on input features, while Filter Attention and Spatial Attention are statically precomputed and depend solely on the parent kernel. The static mechanisms can be efficiently implemented in parallel to reduce the computation cost and parameters of kernel banks (KernelWarehouse [14]) and multi-axis attention (ODConv [15]) by 24%–38%.
@@
where is the base kernel shared across multiple identical convolutional layers, is the modulated kernel for -th layer, and , , and are three components of -th adapter. The represents the dot product operation.
Channel Attention: In fact, applying Channel Attention to convolutional kernels is equivalent to applying it to feature maps. However, for efficient implementation, we adopt the latter approach. Our implementation resembles SE-Net [11] but incorporates a Batch Normalization (BN) layer after the first fully connected (FC) layer. The default channel reduction ratio is set to 4. The detailed implementation process is as follows:
where the denotes the sigmoid function, denotes represents the global average pooling operation.
Spatial Attention and Filter Attention: Compared to the implementation of Channel Attention, Spatial Attention and Filter Attention are simpler. They are implemented as learnable tensors with shapes and , and are stored as static parameters of the adapter. Once trained, these parameters remain fixed. During inference, they can either be fused during initialization or dynamically computed. The Spatial Attention can modulate the spatial distribution of convolutional kernels, enhancing sensitivity to critical regions (e.g., edges or textures). The Filter Attention can adaptively scale output channels, amplifying discriminative features while suppressing noise. So, our proposed Adapter design enables input-aware kernel modulation while retaining the computational benefits of standard convolutions (Fig. 5(b-d)).
@@
Through analysis, we observe that previous works’ purely dynamic computation of spatial attention and filter attention introduces significant computational overhead. By decoupling these components and computing them in parallel, we achieve both efficiency and expressiveness. This decoupled design inspired the name “DNA” for our adapter. The pseudocode implementation is shown in Algorithm 1.

### v2

@@ Abstract
Dynamic convolution enhances model capacity by adaptively combining multiple kernels, yet faces critical trade-offs: prior works either (1) incur significant parameter overhead by scaling kernel numbers linearly, (2) compromise inference speed through complex kernel interactions, or (3) struggle to jointly optimize dynamic attention and static kernels. We observe that pre-trained Convolutional Neural Networks (CNNs) exhibit inter-layer redundancy akin to that in Large Language Models (LLMs). Specifically, dense convolutional layers can be efficiently replaced by derived “child” layers generated from a shared “parent” convolutional kernel through an adapter. To address these limitations and implement the weight-sharing mechanism, we propose a lightweight convolution kernel plug-in, named KernelDNA. It decouples kernel adaptation into input-dependent dynamic routing and pre-trained static modulation, ensuring both parameter efficiency and hardware-friendly inference. Unlike existing dynamic convolutions that expand parameters via multi-kernel ensembles, our method leverages cross-layer weight sharing and adapter-based modulation, enabling dynamic kernel specialization without altering the standard convolution structure. This design preserves the native computational efficiency of standard convolutions while enhancing representation power through input-adaptive kernel adjustments. Experiments on image classification and dense prediction tasks demonstrate that KernelDNA achieves state-of-the-art accuracy-efficiency balance among dynamic convolution variants.
@@ 1 Introduction
Convolutional Neural Networks [he2016deep, sandler2018mobilenetv2, liu2022convnet] have long been built on static convolutional kernels within each layer. While these static kernels enable computationally efficient inference, their fixed, input-agnostic nature inherently limits representation flexibility. Dynamic convolution [Yang2019, Chen2020] attempts to address this by adaptively modulating kernels based on input content, but it inevitably scales parameters linearly with the number of kernels, introducing prohibitive memory costs that hinder real-world deployment. Existing model compression techniques, such as pruning [Li2016], quantization [Esser2019], and knowledge distillation [Gou2021], aim to alleviate parameter growth but require complex multi-stage pipelines: they often depend on fine-tuning pre-trained models with carefully tuned hyperparameters. In contrast, weight-sharing [Wang2024a, Cao2024] strategies in LLMs demonstrate that parameter redundancy can be exploited for drastic compression without performance loss. Inspired by this, we propose integrating dynamic neural networks’ adaptive capacity with LLM-like weight sharing, unlocking the potential of CNNs through a novel parameter-efficient dynamic modulation framework.
@@
To validate the inherent inter-layer redundancy in conventional CNNs, we visualize the similarity between convolutional kernels across layers using the Linear Centered Kernel Alignment (CKA) index [Kornblith2019]. As shown in Figure 1, except for a few layers that exhibit low similarity with others, the majority of layers show high similarity. This observation suggests that compression or transformation can be applied to enable similar layers to share a base layer, thereby reducing the model’s parameters without compromising the architecture. Drawing inspiration from Epigenetics [waddington1942epigenotype]—where phenotypic traits are regulated through non-heritable modifications (e.g., DNA methylation) without altering the underlying genetic code—we design a lightweight adapter module. This module decouples input-dependent dynamic channel attention from kernel-centric spatial modulation, enabling a shared “parent” kernel within a block or stage to generate derived “child” layers via adapter-based transformations. By replacing dense convolutional layers with these adapted child layers, we achieve substantial parameter reduction while preserving performance. Figure 2 illustrates the distinction between the kernel sharing mechanism in our method and mainstream dynamic convolution approaches.
@@
Both previous paradigms (i.e.,Figure 2 (b) and (c)) share critical limitations: they introduce computationally intensive operations (e.g., dynamic kernel assembly or generation) that degrade inference speed and often cause parameter inflation. In contrast, our kernel-sharing strategy not only reduces parameters but also preserves the efficient inference advantage of traditional convolution by reusing and modulating shared base kernels across layers. Notably, parameter sharing introduces implicit regularization [Lan2019], mitigating overfitting risks and even enhancing generalization.
@@
To ensure hardware-friendly efficiency, our adapter integrates three enhanced attention mechanisms: Channel Attention, Filter Attention, and Spatial Attention. These mechanisms allow each convolutional layer to retain distinct characteristics while benefiting from cross-layer captured through shared parameters. Termed KernelDNA, this plug-and-play module functions as an epigenetic “marker”, dynamically tailoring kernel behavior without altering the base parameter structure. It is universally applicable to any standard convolution operator, including regular convolutions, depthwise convolutions, and 11 convolutions.
@@
Extensive experiments on ImageNet-1K and downstream tasks demonstrate that KernelDNA surpasses prior dynamic convolution methods in accuracy while preserving near-original inference speeds (e.g., 7,683 fps vs. ResNet18’s 8,463 fps, 90.8% throughput retention). Compared to SOTA variants like FDConv (11.75M params), KernelDNA achieves 1.27 parameter reduction (9.24M for ResNet18) with higher accuracy (74.23% vs. 71.25% Top-1). On lightweight models like MobileNetV2, KernelDNA attains 75.51% Top-1 accuracy (vs. KernelWarehouse’s 74.68%) using only 3.14M parameters (vs. KernelWarehouse’s 5.17M), while maintaining 99% of the base model’s throughput (6,260 fps vs. 6,323 fps). These results establish a new Pareto frontier, where KernelDNA uniquely balances parameter efficiency, hardware compatibility, and adaptive performance across diverse architectures.
@@ 2 Related Work
Dynamic Convolution. Dynamic convolution aims to enhance the flexibility of CNNs by allowing kernel weights to adapt to input data, breaking the limitations of static convolutions. Early methods like CondConv [Yang2019] use input-dependent routing to combine multiple fixed kernels, while DY-CNNs [Chen2020] leverage channel attention to control dynamic aggregation. Later, approaches such as DCD [Li2021] and DD [Zhou2021] introduce low-rank decompositions or separate channel and spatial modulation, enabling finer adaptive behavior with reduced overhead. More recent designs, including ODConv [Li2022] and SD-Conv [He2023], further generalize adaptiveness to spatial sizes or sparsify kernels for efficiency, but often at the cost of increased complexity, memory, or unstable latency. KernelWarehouse [Li2024] assembles kernels from a large bank of fine-grained components, and FDConv [Chen2025b] adapts kernels in the frequency domain. However, despite improvements, most dynamic convolution methods either introduce heavy memory and latency overhead due to complex kernel synthesis at inference, or trade off accuracy for efficiency. Our KernelDNA addresses these shortcomings by adopting cross-layer weight sharing and lightweight adapters, delivering effective dynamic modulation with minimal inference cost and robust throughput.
Weight Sharing. Weight sharing has been widely investigated in NLP, starting with ALBERT [Lan2019] for parameter-efficient transformers by tying parameters across layers. These ideas inspired similar work in vision models, such as parameter multiplexing in MiniViT [Zhang2022], low-rank decomposition in MPOBERT [Liu2023a], RL-based dynamic layer reuse [Hay2024], and SVD-based basis sharing [Wang2024a]. Further, Relaxed Recursive Transformers [Bae2024] integrate LoRA-like adapters for effective sharing. Despite their success in transformer-based models, the application of deep weight sharing in CNNs, particularly under dynamic convolution settings, remains relatively unexplored. Our work is among the first to systematically integrate weight sharing and dynamic convolution, achieving both high parameter efficiency and fast inference.
Visual Attention Modules. Visual attention mechanisms are now integral to improving CNN feature representations. SE-Net [hu2018squeeze] introduced channel-wise attention via global pooling and fully connected layers, while CBAM [woo2018cbam] further augments channel attention with spatial attention for joint modulation. ECA-Net [wang2020eca] simplifies channel interaction using efficient local 1D convolutions, and CGC [lin2020context] employs context gating to integrate input and global context for kernel modulation. Different from these, KernelDNA separates dynamic channel attention from static filter and spatial attention, employing a lightweight adapter that minimizes computational costs, thereby improving both adaptability and efficiency in practical CNN deployments.
@@ 3 Preliminaries
Standard Convolution. Given an input feature map with channels and spatial dimensions , the standard convolution operation applies a fixed filter to produce an output feature map . At each output spatial location , the convolution is computed as:
where is the output channel index, is the input channel index, are the spatial indices, and is the bias term. The spatial indices are adjusted with appropriate padding to ensure valid indexing. The filter is shared across all spatial locations, making the operation computationally efficient but input-agnostic. Here, the channel dimension is explicitly summed over, ensuring proper alignment between the input channels and the kernel’s input channel dimension.
Conventional Dynamic Convolution. Dynamic convolution improves efficiency by aggregating a set of static kernels where each using input-dependent attention weights :
where is a scalar weight computed via a lightweight attention mechanism such that . The aggregated kernel maintains the same dimensions as each individual kernel and is then applied using standard convolution.
@@ 4.1 Dynamic Kernel Sharing
Recent works have aimed to enhance the flexibility of dynamic convolution by increasing kernel diversity or granularity. For instance, KernelWarehouse [Li2024] decomposes each convolutional kernel into hundreds of fine-grained cell-wise components , building a global bank of reusable kernel elements. During inference, it assembles full kernels on-the-fly using input-driven, contrast-based attention, then applies them to feature maps. FDConv [Chen2025b] instead leverages Fourier-based kernels for improved adaptability but introduces significant memory consumption and additional computational steps. Although KernelWarehouse advertises cross-layer kernel sharing, it still relies on aggregating candidate static kernels via attention—much like classic kernel bank approaches, just at finer granularity (see Figure 3 (a)). And this strategy calculates attention over many components of kernels and synthesizes full kernels at runtime, leading to considerable computational overhead. Such operations demand excessive memory access and are unfriendly to hardware, resulting in much slower inference than standard convolutions and limiting deployment in latency-sensitive applications.
In contrast to KernelWarehouse’s fine-grained decomposition and reassembly, our KernelDNA preserves the structure of standard convolutions while introducing layer-wise parameter sharing. Instead of complex computations, KernelDNA employs lightweight adapters to generate layer-specific kernels from shared base parameters, retaining the inference efficiency of standard convolutions. This design enables seamless replacement of traditional convolutions, as depicted in Figure 3 (b), where the solid box representations a static convolutional kernel, which serves as the parent layer in KernelDNA. The dashed boxes denote child layers derived from the parent layer through adapter transformations. These child layers do not have their own weights, which are dynamically generated during inference to perform convolution operations.
Furthermore, our adapter integrates both static attention (determined during training) and input-dependent dynamic attention (computed at inference), decoupling purely dynamic attention mechanisms of previous works. This dual-pathway design enables flexible adaptation to different deployment scenarios: dynamic attention can be disabled for minimal latency with only a slight accuracy trade-off, or enabled to boost performance at the cost of marginal additional computational overhead. Thus, KernelDNA can serve as a plug-and-play replacement for conventional convolution layers, balancing efficiency and adaptability without compromising hardware compatibility.
@@ 4.2 Light-weight Modulation Adapter
To efficiently transform base (parent) kernels into shared (child) kernels, we design a lightweight adapter that incorporates three attention mechanisms to modulate kernel properties across different dimensions: input channel, output channel, and spatial interactions. These mechanisms are termed Channel Attention, Filter Attention, and Spatial Attention, respectively. As illustrated in Figure 4, Channel Attention is dynamically computed based on input features, while Filter Attention and Spatial Attention are statically stored as model parameters. The static mechanisms can be efficiently implemented in parallel, reducing the computational overhead and parameter count compared to purely dynamic attention mechanisms such as ODConv.
@@
where is the base kernel shared across multiple identical convolutional layers, is the modulated kernel for -th layer, and , , and are three components of -th adapter. The represents element-wise multiplication with broadcasting.
Channel Attention: In fact, applying Channel Attention to convolutional kernels is equivalent to applying it to feature maps. We adopt the latter approach for efficient implementation. Our implementation resembles SE-Net [hu2018squeeze] but incorporates a Batch Normalization (BN) layer after the first fully connected (FC) layer. The default channel reduction ratio is set to 4. The detailed implementation process is as follows:
where the denotes the sigmoid function, denotes the global average pooling operation.
Spatial Attention and Filter Attention: Compared to the implementation of Channel Attention, Spatial Attention and Filter Attention are simpler. They are implemented as learnable tensors with dimensions and , and are stored as static parameters of the adapter. Once trained, these parameters remain fixed and can be fused with the parent kernel during initialization for inference. The Spatial Attention modulates the spatial distribution of convolutional kernels, enhancing sensitivity to critical regions (e.g., edges or textures) [woo2018cbam]. The Filter Attention adaptively scales output channels, amplifying discriminative features while suppressing noise. The design enables input-aware kernel modulation while retaining the computational benefits of standard convolutions (Figure 4(b-d)).
@@
Through analysis, we observe that previous works’ purely dynamic computation of spatial attention and filter attention introduces significant storage overhead. By decoupling these components and applying them in parallel, we achieve both efficiency and expressiveness. This decoupled design inspired the name “DNA” for our adapter. The pseudocode implementation is shown in Algorithm 1.